Test of L^{A}TEX font family ‘txr’

Font sample produced with ‘fontsmpl’

March 10, 2005

Test of font T1/txr/m/n. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘Ð’ \DJ is ‘Ð’ \L is ‘Ł’ \NG is ‘Ŋ’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘Þ’ \_is ‘_’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘ð’ \dj is ‘đ’

\dots is ‘. . . ’ \guillemotleft is ‘«’ \guillemotright is ‘»’
\guilsinglleft is ‘‹’ \guilsinglright is ‘›’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘ł’ \ng is ‘ŋ’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘£’

\quotedblbase is ‘„’ \quotesinglbase is ‘‚’ \ss is ‘ß’
\textasciicircum is ‘^’ \textasciitilde is ‘~’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’

\textquotedbl is ‘"’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘␣’ \th is ‘þ’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A.C.D.E.G.I.L.N.O.R.S.T.U.Y.Z.a.c.d.e.g.ı.i.l.n.o.r.s.t.u.y.z.
\k ĄC̨D̨ĘG̨ĮL̨N̨ǪR̨S̨T̨ŲY̨Z̨ąc̨d̨ęg̨įįl̨n̨ǫr̨s̨t̨ųy̨z̨
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font OT1/txr/m/n. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘?’ \DJ is ‘?’ \L is ‘�’ \NG is ‘?’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘?’ \_is ‘ ’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘?’ \dj is ‘?’

\dots is ‘. . . ’ \guillemotleft is ‘?’ \guillemotright is ‘?’
\guilsinglleft is ‘?’ \guilsinglright is ‘?’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘�’ \ng is ‘?’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘$’

\quotedblbase is ‘?’ \quotesinglbase is ‘?’ \ss is ‘ß’
\textasciicircum is ‘ˆ ’ \textasciitilde is ‘˜ ’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘?’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘ ’ \th is ‘?’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A.C.D.E.G.I.L.N.O.R.S.T.U.Y.Z.a.c.d.e.g.ı.i.l.n.o.r.s.t.u.y.z.
\k ?A?C?D?E?G?I?L?N?O?R?S?T?U?Y?Z?a?c?d?e?g?ı?i?l?n?o?r?s?t?u?y?z
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font T1/txr/m/it. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘Ð’ \DJ is ‘Ð’ \L is ‘Ł’ \NG is ‘Ŋ’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘Þ’ \_is ‘_’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘ð’ \dj is ‘đ’

\dots is ‘. . . ’ \guillemotleft is ‘«’ \guillemotright is ‘»’
\guilsinglleft is ‘‹’ \guilsinglright is ‘›’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘ł’ \ng is ‘ŋ’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘£’

\quotedblbase is ‘„’ \quotesinglbase is ‘‚’ \ss is ‘ß’
\textasciicircum is ‘^’ \textasciitilde is ‘~’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘"’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘␣’ \th is ‘þ’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡī̄il̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A .C .D .E.G .I.L .N .O .R .S.T .U .Y .Z .a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.
\k ĄC̨D̨ĘG̨ĮL̨N̨ǪR̨S̨T̨ŲY̨Z̨ąc̨d̨ęg̨įįl̨n̨ǫr̨s̨t̨ųy̨z̨
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font OT1/txr/m/it. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy s oufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘?’ \DJ is ‘?’ \L is ‘�’ \NG is ‘?’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘?’ \_is ‘ ’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘?’ \dj is ‘?’

\dots is ‘. . . ’ \guillemotleft is ‘?’ \guillemotright is ‘?’
\guilsinglleft is ‘?’ \guilsinglright is ‘?’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘�’ \ng is ‘?’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘$’

\quotedblbase is ‘?’ \quotesinglbase is ‘?’ \ss is ‘ß’
\textasciicircum is ‘ˆ ’ \textasciitilde is ‘˜ ’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘?’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘ ’ \th is ‘?’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡī̄il̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A .C .D .E.G .I.L .N .O .R .S.T .U .Y .Z .a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.
\k ?A?C?D?E?G?I?L?N?O?R?S?T?U?Y?Z?a?c?d?e?g?ı?i?l?n?o?r?s?t?u?y?z
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font T1/txr/m/sl. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘Ð’ \DJ is ‘Ð’ \L is ‘Ł’ \NG is ‘Ŋ’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘Þ’ \_is ‘_’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘ð’ \dj is ‘đ’

\dots is ‘. . . ’ \guillemotleft is ‘«’ \guillemotright is ‘»’
\guilsinglleft is ‘‹’ \guilsinglright is ‘›’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘ł’ \ng is ‘ŋ’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘£’

\quotedblbase is ‘„’ \quotesinglbase is ‘‚’ \ss is ‘ß’
\textasciicircum is ‘^’ \textasciitilde is ‘~’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘"’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘␣’ \th is ‘þ’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A .C.D .E.G .I.L.N .O .R.S.T.U .Y .Z.a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.

\k ĄC̨D̨ĘG̨ĮL̨N̨ǪR̨S̨T̨ŲY̨Z̨ąc̨d̨ęg̨įįl̨n̨ǫr̨s̨t̨ųy̨z̨
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font OT1/txr/m/sl. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘?’ \DJ is ‘?’ \L is ‘�’ \NG is ‘?’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘?’ \_is ‘ ’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘?’ \dj is ‘?’

\dots is ‘. . . ’ \guillemotleft is ‘?’ \guillemotright is ‘?’
\guilsinglleft is ‘?’ \guilsinglright is ‘?’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘�’ \ng is ‘?’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘$’

\quotedblbase is ‘?’ \quotesinglbase is ‘?’ \ss is ‘ß’
\textasciicircum is ‘ˆ ’ \textasciitilde is ‘˜ ’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘?’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘ ’ \th is ‘?’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź

\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A .C .D .E .G .I.L.N .O .R .S.T .U .Y .Z.a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.
\k ?A?C?D?E?G?I?L?N?O?R?S?T?U?Y?Z?a?c?d?e?g?ı?i?l?n?o?r?s?t?u?y?z
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font T1/txr/m/sc. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘Ð’ \DJ is ‘Ð’ \L is ‘Ł’ \NG is ‘Ŋ’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘Þ’ \_is ‘_’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘ð’ \dj is ‘đ’

\dots is ‘. . . ’ \guillemotleft is ‘«’ \guillemotright is ‘»’
\guilsinglleft is ‘‹’ \guilsinglright is ‘›’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘ł’ \ng is ‘ŋ’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘£’

\quotedblbase is ‘„’ \quotesinglbase is ‘‚’ \ss is ‘ß’
\textasciicircum is ‘^’ \textasciitilde is ‘~’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’

\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘"’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘␣’ \th is ‘þ’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A.C.D.E.G.I.L.N.O.R.S.T.U.Y.Z.a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.
\k ĄC̨D̨ĘG̨ĮL̨N̨ǪR̨S̨T̨ŲY̨Z̨ąc̨d̨ęg̨įįl̨n̨ǫr̨s̨t̨ųy̨z̨
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font OT1/txr/m/sc. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (ffiTHE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) fflBut aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘?’ \DJ is ‘?’ \L is ‘�’ \NG is ‘?’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘?’ \_is ‘ ’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘?’ \dj is ‘?’

\dots is ‘. . . ’ \guillemotleft is ‘?’ \guillemotright is ‘?’
\guilsinglleft is ‘?’ \guilsinglright is ‘?’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘�’ \ng is ‘?’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘$’

\quotedblbase is ‘?’ \quotesinglbase is ‘?’ \ss is ‘ß’
\textasciicircum is ‘ˆ ’ \textasciitilde is ‘˜ ’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘?’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘ ’ \th is ‘?’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A .C .D .E .G .I.L.N .O .R .S.T .U .Y .Z.a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.
\k ?A?C?D?E?G?I?L?N?O?R?S?T?U?Y?Z?a?c?d?e?g?ı?i?l?n?o?r?s?t?u?y?z
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font T1/txr/bx/n. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.”

(¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘Ð’ \DJ is ‘Ð’ \L is ‘Ł’ \NG is ‘Ŋ’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘Þ’ \_is ‘_’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘ð’ \dj is ‘đ’

\dots is ‘. . . ’ \guillemotleft is ‘«’ \guillemotright is ‘»’
\guilsinglleft is ‘‹’ \guilsinglright is ‘›’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘ł’ \ng is ‘ŋ’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘£’

\quotedblbase is ‘„’ \quotesinglbase is ‘‚’ \ss is ‘ß’
\textasciicircum is ‘^’ \textasciitilde is ‘~’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘"’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘␣’ \th is ‘þ’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A.C.D.E.G.I.L.N.O.R.S.T.U.Y.Z.a.c.d.e.g.ı.i.l.n.o.r.s.t.u.y.z.
\k ĄC̨D̨ĘG̨ĮL̨N̨ǪR̨S̨T̨ŲY̨Z̨ąc̨d̨ęg̨įįl̨n̨ǫr̨s̨t̨ųy̨z̨
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font OT1/txr/bx/n. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘?’ \DJ is ‘?’ \L is ‘� ’ \NG is ‘?’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘?’ \_is ‘ ’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘?’ \dj is ‘?’

\dots is ‘. . . ’ \guillemotleft is ‘?’ \guillemotright is ‘?’
\guilsinglleft is ‘?’ \guilsinglright is ‘?’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘�’ \ng is ‘?’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘$’

\quotedblbase is ‘?’ \quotesinglbase is ‘?’ \ss is ‘ß’
\textasciicircum is ‘ˆ ’ \textasciitilde is ‘˜ ’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘?’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘ ’ \th is ‘?’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A.C.D.E.G.I.L.N.O.R.S.T.U.Y.Z.a.c.d.e.g.ı.i.l.n.o.r.s.t.u.y.z.

\k ?A?C?D?E?G?I?L?N?O?R?S?T?U?Y?Z?a?c?d?e?g?ı?i?l?n?o?r?s?t?u?y?z
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font T1/txr/bx/it. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘Ð’ \DJ is ‘Ð’ \L is ‘Ł’ \NG is ‘Ŋ’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘Þ’ \_is ‘_’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘ð’ \dj is ‘đ’

\dots is ‘. . . ’ \guillemotleft is ‘«’ \guillemotright is ‘»’
\guilsinglleft is ‘‹’ \guilsinglright is ‘›’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘ł’ \ng is ‘ŋ’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘£’

\quotedblbase is ‘„’ \quotesinglbase is ‘‚’ \ss is ‘ß’
\textasciicircum is ‘^’ \textasciitilde is ‘~’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘"’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘␣’ \th is ‘þ’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈

\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡī̄īln̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A .C .D .E.G .I.L.N .O .R .S.T.U .Y .Z.a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.
\k ĄC̨D̨ĘG̨ĮL̨N̨ǪR̨S̨T̨ŲY̨Z̨ąc̨d̨ęg̨įįl̨n̨ǫr̨s̨t̨ųy̨z̨
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font OT1/txr/bx/it. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘?’ \DJ is ‘?’ \L is ‘�’ \NG is ‘?’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘?’ \_is ‘ ’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘?’ \dj is ‘?’

\dots is ‘. . . ’ \guillemotleft is ‘?’ \guillemotright is ‘?’
\guilsinglleft is ‘?’ \guilsinglright is ‘?’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘�’ \ng is ‘?’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘$’

\quotedblbase is ‘?’ \quotesinglbase is ‘?’ \ss is ‘ß’
\textasciicircum is ‘ˆ ’ \textasciitilde is ‘˜ ’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’

\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘?’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘ ’ \th is ‘?’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡī̄īln̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A .C .D .E.G .I.L.N .O .R .S.T.U .Y .Z.a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.
\k ?A?C?D?E?G?I?L?N?O?R?S?T?U?Y?Z?a?c?d?e?g?ı?i?l?n?o?r?s?t?u?y?z
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font T1/txr/bx/sl. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘Ð’ \DJ is ‘Ð’ \L is ‘Ł’ \NG is ‘Ŋ’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘Þ’ \_is ‘_’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘ð’ \dj is ‘đ’

\dots is ‘. . . ’ \guillemotleft is ‘«’ \guillemotright is ‘»’
\guilsinglleft is ‘‹’ \guilsinglright is ‘›’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘ł’ \ng is ‘ŋ’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘£’

\quotedblbase is ‘„’ \quotesinglbase is ‘‚’ \ss is ‘ß’
\textasciicircum is ‘^’ \textasciitilde is ‘~’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘"’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘␣’ \th is ‘þ’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A .C.D .E.G .I.L.N .O .R.S.T.U .Y .Z.a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.
\k ĄC̨D̨ĘG̨ĮL̨N̨ǪR̨S̨T̨ŲY̨Z̨ąc̨d̨ęg̨įįl̨n̨ǫr̨s̨t̨ųy̨z̨
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font OT1/txr/bx/sl. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.”

(¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘?’ \DJ is ‘?’ \L is ‘� ’ \NG is ‘?’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘?’ \_is ‘ ’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘?’ \dj is ‘?’

\dots is ‘. . . ’ \guillemotleft is ‘?’ \guillemotright is ‘?’
\guilsinglleft is ‘?’ \guilsinglright is ‘?’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘�’ \ng is ‘?’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘$’

\quotedblbase is ‘?’ \quotesinglbase is ‘?’ \ss is ‘ß’
\textasciicircum is ‘ˆ ’ \textasciitilde is ‘˜ ’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘?’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘ ’ \th is ‘?’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A .C.D .E.G.I.L.N .O.R.S.T.U .Y .Z.a.c.d.e.g.ı .i.l.n.o.r.s.t.u.y.z.
\k ?A?C?D?E?G?I?L?N?O?R?S?T?U?Y?Z?a?c?d?e?g?ı?i?l?n?o?r?s?t?u?y?z
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font T1/txr/bx/sc. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (¡THE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) ¿But aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘Ð’ \DJ is ‘Ð’ \L is ‘Ł’ \NG is ‘Ŋ’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘Þ’ \_is ‘_’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘ð’ \dj is ‘đ’

\dots is ‘. . . ’ \guillemotleft is ‘«’ \guillemotright is ‘»’
\guilsinglleft is ‘‹’ \guilsinglright is ‘›’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘ł’ \ng is ‘ŋ’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘£’

\quotedblbase is ‘„’ \quotesinglbase is ‘‚’ \ss is ‘ß’
\textasciicircum is ‘^’ \textasciitilde is ‘~’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘"’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘␣’ \th is ‘þ’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A.C.D.E.G.I.L.N.O.R.S.T.U.Y.Z.a.c.d.e.g.ı.i.l.n.o.r.s.t.u.y.z.

\k ĄC̨D̨ĘG̨ĮL̨N̨ǪR̨S̨T̨ŲY̨Z̨ąc̨d̨ęg̨įįl̨n̨ǫr̨s̨t̨ųy̨z̨
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃

Test of font OT1/txr/bx/sc. Some text:

On November 14, 1885, Senator & Mrs. Leland Stanford called together at their San Francisco mansion the 24 prominent men who had been chosen as the first trustees of The Leland Stanford Junior University. They handed to the board the Founding Grant of the University, which they had executed three days before. This document—with various amendments, legislative acts, and court decrees—remains as the University’s charter. In bold, sweeping language it stipulates that the objectives of the University are “to qualify students for personal success and direct usefulness in life; and to promote the publick welfare by exercising an influence in behalf of humanity and civilization, teaching the blessings of liberty regulated by law, and inculcating love and reverence for the great principles of government as derived from the inalienable rights of man to life, liberty, and the pursuit of happiness.” (ffiTHE DAZED BROWN FOX QUICKLY GAVE 12345–67890 JUMPS!) fflBut aren’t Kafka’s Schloß and Æsop’s Œuvres often naïve vis-à-vis the dæmonic phœnix’s official rôle in fluffy soufflés?

Some glyphs:

\# is ‘#’ \$ is ‘$’ \% is ‘%’ \& is ‘&’ \AA is ‘Å’ \AE is ‘Æ’ \DH is ‘?’ \DJ is ‘?’ \L is ‘� ’ \NG is ‘?’ \OE is ‘Œ’ \O is ‘Ø’ \P is ‘¶’ \SS is ‘SS’ \S is ‘§’ \TH is ‘?’ \_is ‘ ’ \aa is ‘å’ \ae is ‘æ’ \copyright is ‘\begin{matrix}c \\ ◯ \end{matrix} ’ \dag is ‘†’ \ddag is ‘‡’ \dh is ‘?’ \dj is ‘?’

\dots is ‘. . . ’ \guillemotleft is ‘?’ \guillemotright is ‘?’
\guilsinglleft is ‘?’ \guilsinglright is ‘?’ \i is ‘ı’ \j is ‘ȷ’

\l is ‘�’ \ng is ‘?’ \oe is ‘œ’ \o is ‘ø’ \pounds is ‘$’

\quotedblbase is ‘?’ \quotesinglbase is ‘?’ \ss is ‘ß’
\textasciicircum is ‘ˆ ’ \textasciitilde is ‘˜ ’
\textbackslash is ‘\’ \textbar is ‘|’ \textbullet is ‘•’
\textcompwordmark is ‘’ \textemdash is ‘—’ \textendash is ‘–’
\textexclamdown is ‘¡’ \textgreater is ‘>’
\texthyphenchar is undefined \textless is ‘<’
\textperiodcentered is ‘·’ \textquestiondown is ‘¿’
\textquotedblleft is ‘“’ \textquotedblright is ‘”’
\textquotedbl is ‘?’ \textquoteleft is “’ \textquoteright is ‘”
\textvisiblespace is ‘ ’ \th is ‘?’ \{ is ‘{’ \} is ‘}’

Some accents:

\" ÄC̈D̈ËG̈ÏL̈N̈ÖR̈S̈T̈ÜŸZ̈äc̈d̈ëg̈ïïl̈n̈ör̈s̈ẗüÿz̈
\' ÁĆD́ÉǴÍĹŃÓŔŚT́ÚÝŹáćd́éǵííĺńóŕśt́úýź
\. ȦĊḊĖĠİL̇ṄȮṘṠṪU̇ẎŻȧċḋėġiil̇ṅȯṙṡṫu̇ẏż
\= ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\H A̋C̋D̋E̋G̋I̋L̋N̋ŐR̋S̋T̋ŰY̋Z̋a̋c̋d̋e̋g̋i̋i̋l̋n̋őr̋s̋t̋űy̋z̋
\^ÂĈD̂ÊĜÎL̂N̂ÔR̂ŜT̂ÛŶẐâĉd̂êĝîîl̂n̂ôr̂ŝt̂ûŷẑ
\` ÀC̀D̀ÈG̀ÌL̀ǸÒR̀S̀T̀ÙỲZ̀àc̀d̀èg̀ììl̀ǹòr̀s̀t̀ùỳz̀
\b ĀC̄D̄ĒḠĪL̄N̄ŌR̄S̄T̄ŪȲZ̄āc̄d̄ēḡīīl̄n̄ōr̄s̄t̄ūȳz̄
\c A̧ÇḐȨĢI̧ĻŅO̧ŖŞŢU̧Y̧Z̧a̧çḑȩģi̧i̧ļņo̧ŗşţu̧y̧z̧
\d A.C.D.E.G.I.L.N.O.R.S.T.U.Y.Z.a.c.d.e.g.ı.i.l.n.o.r.s.t.u.y.z.
\k ?A?C?D?E?G?I?L?N?O?R?S?T?U?Y?Z?a?c?d?e?g?ı?i?l?n?o?r?s?t?u?y?z
\u ĂC̆D̆ĔĞĬL̆N̆ŎR̆S̆T̆ŬY̆Z̆ăc̆d̆ĕğĭĭl̆n̆ŏr̆s̆t̆ŭy̆z̆
\v ǍČĎĚǦǏĽŇǑŘŠŤǓY̌Žǎčďěǧǐǐľňǒřšťǔy̌ž
\~ ÃC̃D̃ẼG̃ĨL̃ÑÕR̃S̃T̃ŨỸZ̃ãc̃d̃ẽg̃ĩĩl̃ñõr̃s̃t̃ũỹz̃